### v4

# Reinforced Generation of Combinatorial Structures: Applications to Complexity Theory

Abstract
We explore whether techniques from AI can help discover new combinatorial structures that improve on known limits on efficient algorithms. Specifically, we use AlphaEvolve (an LLM coding agent) to study two settings:
a) Average-case hardness for MAX-CUT and MAX-Independent Set: We improve a recent result of Kunisky and Yu to obtain near-optimal upper and (conditional) lower bounds on certification algorithms for MAX-CUT and MAX-Independent Set on random 3- and 4-regular graphs. Our improved lower bounds are obtained by constructing nearly extremal Ramanujan graphs on as many as $163$ nodes, using AlphaEvolve. Additionally, via analytical arguments we strengthen the upper bounds to settle the computational hardness of these questions up to an error in the third decimal place.
b) Worst-case Hardness of Approximation for MAX-k-CUT: We obtain new inapproximability results, proving that it is NP-hard to approximate $\textup{MAX-}{4}\textup{-CUT}$ and $\textup{MAX-}{3}\textup{-CUT}$ within factors of $0.987$ and $0.9649$ respectively, using AlphaEvolve to discover new gadget reductions. Our $\textup{MAX-}{4}\textup{-CUT}$ result improves upon the SOTA of $0.9883$, and our $\textup{MAX-}{3}\textup{-CUT}$ result improves on the current best gadget-based inapproximability result of $0.9853$, but falls short of improving the SOTA of $16/17$ that relies on a custom PCP, rather than a gadget reduction from “standard” Håstad-style PCPs.
A key technical challenge we faced: verifying a candidate construction produced by AlphaEvolve is costly (often requiring exponential time). In both settings above, our results were enabled by using AlphaEvolve itself to evolve the verification procedure to be faster (sometimes by $10,000\times$). We conclude with a discussion of norms by which to assess the assistance from AI in developing proofs.
1 Introduction
In this paper, we address the following question: Can AI based methods help us make novel and non-trivial discoveries in complexity theory? We provide evidence towards answering this in the affirmative, via progress on two problems.
  • a)
    Hardness of certifying upper bounds on the MAX-CUT and MAX-IND-SET of sparse random graphs: We improve upon the combinatorial structures presented in [26], achieving improved lower bounds in the $\{3,4\}$-regular cases. Together with new non-AI based algorithmic improvements [^1] These algorithmic improvements are derived from a discussion with Sidhanth Mohanty, Northwestern University., we achieve an almost tight resolution of these questions [^2] By almost tight, we mean that the lower bounds and upper bounds on efficient certificates match up to the second decimal place..
  • b)
    NP-Hardness of approximation for $\textup{MAX-}{k}\textup{-CUT}$. We find new gadget-based reductions (from standard Håstad PCPs [19]) for the NP-hardness of approximating $\textup{MAX-}{k}\textup{-CUT}$ for $k\in\{3,4\}$. For $\textup{MAX-}{4}\textup{-CUT}$, we improve on the current best inapproximability result from $0.9883$ [2] to $0.987$. For $\textup{MAX-}{3}\textup{-CUT}$, we improve on the current best gadget-based inapproximability result from $0.9853$ [24] to $0.9649$. Our result for $\textup{MAX-}{3}\textup{-CUT}$ is sandwiched between two results using custom PCPs for this problem: [14] with $0.9696$, and [2] with $0.9411$.
To this end we invoke AlphaEvolve [30], a coding agent for scientific and algorithmic discovery, to find and verify finite combinatorial structures that improve on prior constructions. At a high level AlphaEvolve is a tool that uses an LLM (Large-Language-Model) to iteratively evolve code snippets that generate combinatorial structures (we will sometimes refer to these as constructions) of high quality by some criterion. It is worth emphasizing that while the structures we generate using AlphaEvolve are finite constructions, our main results (Theorems 3.1 and 4.1) embed the universal quantifier $\forall n$, with $n$ being the size of the problem instance, thanks to appropriate “lifting” [26, 36] arguments.
Operationally, the system consists of three parts: a) a code snippet ($\mathbf{C}$) that constructs combinatorial structures, b) An evaluation function (referred to as the verifier) that verifies and scores the generated structures, and c) An LLM that suggests a new search function $\mathbf{C}_{\texttt{new}}$ based on the set of prior $\mathbf{C}$’s and the history of constructions. Via prompting the LLM, the objective is to nudge $\mathbf{C}_{\texttt{new}}$ towards better structures. (Section 2 provides a more detailed overview of AlphaEvolve.)
The efficacy of AlphaEvolve crucially depends on the verification step which in turn conditions the search landscape. Because we seek extremal structures in a large search space, fast verification (including scoring) of a construction helps AlphaEvolve try a large number of combinatorial structures, and learn patterns from them to prune the search space. The eventual sizes of the structures we find for our problems are directly correlated with the speed of verification.
Speeding up verification in AlphaEvolve: The constructions we search for in both problems are undirected graphs. For the hardness of certifying upper bounds for sparse random graphs, we use AlphaEvolve to search for $d=\{3,4\}$-regular Ramanujan graphs, along with a witness for a large cut/independent set. Evaluating whether a graph is Ramanujan, and whether or not the cut/independent set is a valid witness, are efficient. As a result, for this problem we could find and verify graphs with as many as $163$ nodes.
However in studying the NP-Hardness of approximation for $\textup{MAX-}{k}\textup{-CUT}$, we seek gadgets that reduce the 3LIN(k) problem to MAX-k-CUT; here the verification is unfortunately not efficient: it requires checking the gadget against $\Omega(k^{m})$ linear constraints, where $m$ is the total number of variables in the target problem. A standard methodology in gadget-based hardness reductions is to add auxiliary variables to have flexibility in mapping the source and the target problems [36]. This exacerbates the exponential running time of the verifier (for reference, our final gadgets contain $m=14$ and $m=19$ variables for $k=3$ and $k=4$ respectively). To address this we invoke AlphaEvolve to speed up the verifier, as follows. (In Section 4.1 we discuss this in more detail.)
We began with a code snippet containing a brute-force implementation of the verifier, using AlphaEvolve to iteratively improve its runtime. The main challenge we faced was ensuring that the evolved snippets were correct, in the sense of being functionally equivalent to the original verifier. We found that the following two ideas sufficed to achieve large speedups while maintaining correctness.
  • We used prior code snippets $\mathbf{C}_{\texttt{old}}$, whose correctness has already been verified, to evaluate synthetically generated gadgets. We compare the outputs of a new candidate $\mathbf{C}_{\texttt{new}}$ to this dataset to ensure that it is also correct.
  • We used a separate LLM as a judge to ensure that all candidates $\mathbf{C}_{\texttt{new}}$ are correct implementations of the verifier.
We prompted AlphaEvolve to improve the execution time of the verifier, while ensuring that it passes the two correctness tests above. To our surprise, AlphaEvolve improved the running time by $10,000\times$ for our final $\textup{MAX-}{4}\textup{-CUT}$ gadget of size $m=19$, and by a similar factor for our final $\textup{MAX-}{3}\textup{-CUT}$ gadget of size $m=14$. Furthermore, we found the improved versions to be provably correct by human inspection (given there was a possibility of being correct only on the synthetic dataset). The improvement came from a number of system-level improvements together with a branch-and-bound strategy to prune many of the $\Omega(k^{m})$ constraints.
We emphasize that our theorems do not depend on the correctness of these improved verifiers; even though our AlphaEvolve computations internally use them, the gadgets found by AlphaEvolve were verified using a brute-force algorithm that explicitly checks all $\Omega(k^{m})$ constraints.
2 AlphaEvolve as a Framework for Combinatorial Discovery
AlphaEvolve [30, 34] is a code-mutation LLM agent capable of autonomously evolving a code snippet to hill-climb on a space of objects with a well defined evaluation metric. We use AlphaEvolve, with problem-specific verifiers that score snippets generated by AlphaEvolve, to discover novel finite combinatorial structures that improve results in complexity theory [24, 14, 26]. Previously AlphaEvolve has been used in a variety of scientific domains [34, 30]. In the following, we provide a self-contained description of AlphaEvolve and the verifier. Additionally, we highlight the system enhancements needed to achieve the combinatorial structures in this paper.
Figure 1: Propose-test-refine (PTR) paradigm: Defining combinatorial search with AlphaEvolve. The solid lines define the control flow, and the dashed lines define comments.
2.1 Propose-test-refine (PTR) paradigm
We operate in the PTR paradigm  [^3] The name is motivated from the Propose-test-release (PTR) framework in the differential privacy literature [10], which performs a similar task of testing a proposed candidate construction for (privacy) properties, before releasing it. for setting up experiments to discover novel combinatorial structures. While the paradigm is implicit in prior work [30], we first make it explicit, and then instantiate the extensions needed in this paper. The PTR paradigm consists of three components:
  1. 1.
    Suitably defined conjecture/claim: The first step towards discovery is a concrete definition of the combinatorial structure being sought. For example, it can be a weighted graph with some fixed set of nodes, or a SAT formula with a fixed number of clauses and variables. In this paper, we restrict these to finite structures rather than a parameterized family  [^4] Let $\mathcal{S}$ be a class of combinatorial structures (e.g., graphs, hypergraphs, codes, formulas). A parameterized family of structures from $\mathcal{S}$ is a sequence $\{S_{n}\}_{n\in\mathbb{N}}$ s.t. each $S_{n}\in\mathcal{S}$ is a finite structure whose size depends on the parameter $n$.. While the source and the target problems form a parameterized family (which enables us to prove theorems with $\forall n$ quantification), the gadgets used to translate between them are finite structures. (A gadget may be thought of as a combinatorial structure mapping a source constraint to a linear combination of target constraints [19, 36].)
  2. 2.
    An initial candidate search algorithm: AlphaEvolve operates by making modifications to a code-snippet by asking an LLM to make syntactically valid modifications, while improving some fitness score. Denote the code snippet at time $t$ by $\mathbf{C}_{t}:\mathcal{D}\to\mathcal{S}$, where $\mathcal{D}$ is the domain capturing problem specific parameters and any inputs to the snippet, and $\mathcal{S}$ is the range space of all possible combinatorial structures. Additionally, let $\texttt{score}:\mathcal{S}\to\mathbb{R}$ be the function that assigns a fitness score to a structure. The next function $\mathbf{C}_{t+1}$ is chosen as follows:
    $\mathbf{C}_{t+1}\leftarrow\texttt{AlphaEvolve}\left(\left(\mathbf{C}_{0}(d_{0}),\texttt{score}(\mathbf{C}_{0}(d_{0}))\right),\ldots,\left(\mathbf{C}_{t}(d_{t}),\texttt{score}(\mathbf{C}_{t}(d_{t}))\right)\right)$ (1)
    AlphaEvolve makes an LLM call using the tuple of inputs in the prompt to the LLM, prompting it to output the description of $\mathbf{C}_{t+1}$, with the goal of increasing the score [^5] Technically, AlphaEvolve does not feed in all the prior functions as the prompt to the LLM. Instead, it uses a genetic algorithm to select a subset of them while maintaining high scores, and diversity [34, Figure 1].. We emphasize the following subtle point here: AlphaEvolve does not directly search for combinatorial structures, but instead  a code-snippet that would output a combinatorial structure with a high score. Figure 2 shows an initial code-snippet to be evolved by AlphaEvolve.
    # EVOLVE-BLOCK-START
    def gadget_construction() -> nx.graph:
    """Returns a candidate gadget construction"""
    gadget = nx.Graph()
    # Create a complete graph with random weights
    max_scoring_gadget = gadget
    max_score = evaluate_internal(gadget)
    while time.time() - start_time < 1000.0:
    # Make random modifications to gadget, to create
    # a new gadget: new_gadget
    current_score = evaluate_internal(new_gadget)
    if max_score < current_score:
    max_scoring_gadget = new_gadget
    max_score = current_score
    return max_scoring_gadget
    def evaluate_internal(gadget):
    # Compute the score on the gadget, and return
    return score(gadget)
    # EVOLVE-BLOCK-END
    Figure 2: Initial code-snippet for AlphaEvolve to evolve.
  3. 3.
    A verifier with a smooth scoring function: In the problems we consider, we seek combinatorial structures in spaces of size exponential in the description of the structure. As a result, a well-defined verifier that validates (and scores) the constructions generated by the $\mathbf{C}_{t}$’s (in Equation 1) is necessary. In particular, the verifier serves as a constraint in the optimization landscape.
    Concretely, for average-case hardness in Section 3, this corresponds to validating whether the graphs are Ramanujan with certain properties (e.g., having a large cut, or independent set), and for worst-case NP-hardness in Section 4, this corresponds to verifying whether the soundness and completeness constraints (e.g., [36, 19]) corresponding to the gadget reduction are satisfied.
    As mentioned earlier, every structure generated by AlphaEvolve needs to be scored, including invalid ones that violate certain constraints. The scoring function $\texttt{score}:\mathcal{S}\to\mathbb{R}$ is used to hill-climb on the optimization landscape. For invalid structures, it decides how far from the constraint boundary the current structure lies. (A scoring function is smooth if its value gracefully increases/decreases as one moves away from the constraint boundary.) This is where the most creativity, and problem specific domain expertise comes in. In particular, the efficacy of the scoring function largely decides the overall capability of the system to discover novel structures. Also, scoring is usually the most compute intensive step in the PTR paradigm.
    Our contributions: Our main contributions to the PTR paradigm are the following: a) using AlphaEvolve itself to accelerate the verification (sometimes up to $10,000\times$), and b) an approach towards using AlphaEvolve for solving optimization problems that involve mixed-integer programming (MIP) style constraints. We discuss each of these contributions in detail in Sections 4.1 and 4.2 in the context of gadget discovery for NP-Hardness of approximation.
3 Average Case Hardness for Random Graphs
A central problem in average case complexity is that of certification, where our goal is to certify a property of inputs from some distribution $\mu=\mu_{n}$ over a sample space $\Omega=\Omega_{n}$. In particular, let $P:\Omega\to\{0,1\}$ be a property such that $\mathop{\mathbf{Pr}}_{x\sim\mu}[P(x)=1]=1-o_{n}(1)$. We are given as input a sample $x\sim\mu$, and the goal is to efficiently certify that $x$ satisfies $P$ with high probability.
Definition 3.1.
An efficient certificate of $P$ is any property $P^{\prime}:\Omega\to\{0,1\}$ that is computable in polynomial time such that the following conditions hold.
  1. 1.
    $P^{\prime}(G)=1$ implies $P(G)=1$.
  2. 2.
    $\mathop{\mathbf{Pr}}_{G\sim G(n,d)}[P^{\prime}(G)=1]=1-o_{n}(1)$.
We focus specifically on the setting of refutation, where $P$ is a co-NP property. This setting has seen considerable activity, leading to innovation in various algorithmic techniques such as spectral and sum-of-squares algorithms, along with the development of a variety of methods to establish hardness (for instance, [4, 35, 3, 26, 25]).
In particular, we will consider the problems of certifying properties of sparse random graphs, where $\mu=G(n,d)$ is the uniform distribution over random $d$-regular graphs for some fixed constant $d\in\mathbb{N}$. The main question we consider is the following: what is the strongest upper bound one can efficiently certify on the MAX-CUT (or MAX-IND-SET) of a random $d$-regular graph?
Definition 3.2 (Max-Cut fraction, Max-Independent Set fraction).
For an undirected graph $G$, $\textup{MC}(G)$ is defined as the fraction of edges in a maximum cut of $G$. Similarly, $\textup{IS}(G)$ is defined as the fraction of vertices in a maximum independent set of $G$.
Definition 3.3.
$\sigma^{\textup{MC}}_{d}$ is the infimum over all $\sigma$ such that the property $\mathbf{1}\{\textup{MC}(G)\leq\sigma\}$ has an efficient certificate. Similarly, $\sigma^{\textup{IS}}_{d}$ is the infimum over all $\sigma$ such that $\mathbf{1}\{\textup{IS}(G)\leq\sigma\}$ has an efficient certificate.
There is evidence that this is a nontrivial problem exhibiting a certification gap, in the sense that $\textup{MC}(G)$ concentrates around a constant value strictly smaller than $\sigma^{\textup{MC}}_{d}$ for $G\sim G(n,d)$ [3, 26]. So in some sense, these are the simplest refutation-style properties that exhibit interesting complexity-theoretic phenomena. In this work, we focus on accurately characterizing the value of $\sigma^{\textup{MC}}_{d}$ and $\sigma^{\textup{IS}}_{d}$. We remark that the asymptotics of these quantities as $d\to\infty$ are already understood (thanks to [3, 26]). In service of a precise understanding of this problem in all regimes, we focus on the least understood case where $d$ is small; specifically we study $d\in\{3,4\}$.
Towards this goal, Kunisky and Yu introduced a conjecture [^6] We refer the reader to [26] for supporting evidence and intuition. that relates $\sigma^{\textup{MC}}_{d}$ and $\sigma^{\textup{IS}}_{d}$ to a fundamental question in spectral graph theory, specifically about Ramanujan graphs, which are (in a spectral sense) the most “random” deterministic regular graphs.
Definition 3.4 (Ramanujan graphs).
Let $G$ be a $d$-regular multigraph on $n$ vertices with adjacency matrix $A$. Suppose the eigenvalues of $A$ are $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$. We say that $G$ is Ramanujan if $\max_{i>1}|\lambda_{i}|\leq 2\sqrt{d-1}$.
Definition 3.5.
Define $\gamma^{\textup{MC}}_{d}$ (resp. $\gamma^{\textup{IS}}_{d}$) as the supremum of $\textup{MC}(G)$ (resp. $\textup{IS}(G)$) over all $d$-regular Ramanujan graphs $G$.
$\gamma^{\textup{MC}}_{d}$ (resp. $\gamma^{\textup{IS}}_{d}$) can be thought of as the performance of the optimal “spectral” certificate for bounds on MAX-CUT (resp. MAX-IND-SET).
Conjecture 3.1 ([26]).
$\sigma^{\textup{MC}}_{d}\geq\gamma^{\textup{MC}}_{d}$ and $\sigma^{\textup{IS}}_{d}\geq\gamma^{\textup{IS}}_{d}$ for all $d\geq 3$.
In other words, the existence of Ramanujan graphs with notably large MAX-CUT and MAX-IND-SET values implies computational hardness of our certification problem. In particular, Kunisky and Yu [26] exhibited finite $d$-regular Ramanujan graphs for $d\in\{3,4\}$, yielding concrete lower bounds on $\sigma_{d}^{\textup{MC}}$ and $\sigma_{d}^{\textup{IS}}$ that are strong enough to prove a certification gap. On the other hand, there was a large gap between the best known upper bound [23, 18] and their lower bounds (see Table 1 for a comparison). Our results make progress towards a more precise understanding of this computational question. Firstly, we use AlphaEvolve to construct better $d$-regular Ramanujan graphs, exhibiting stronger lower bounds.
Theorem 3.1 (Lower bound).
We have the lower bounds $\gamma_{4}^{\textup{MC}}\geq 113/124$, $\gamma_{3}^{\textup{IS}}\geq 17/36$, and $\gamma_{4}^{\textup{IS}}\geq 74/163$. Under 3.1, this implies the following computational hardness results: $\sigma_{4}^{\textup{MC}}\geq 113/124>0.911$, $\sigma_{3}^{\textup{IS}}\geq 17/36>0.472$, and $\sigma_{4}^{\textup{IS}}\geq 74/163>0.453$.
This Theorem improves upon the results of [26], who prove $\gamma_{4}^{\textup{MC}}\geq 7/8$, $\gamma_{3}^{\textup{IS}}\geq 11/24$, and $\gamma_{4}^{\textup{IS}}\geq 3/7$. An explicit description of the constructions found by AlphaEvolve can be found in Appendix A.1, and a figure of the construction achieving the $\gamma_{4}^{\textup{MC}}$ bound is presented in Figure 3. We complement these hardness results with an algorithmic improvement over Hoffman’s classic bound [23, 18].
Theorem 3.2 (Upper bound).
We have the upper bounds $\sigma_{3}^{\textup{MC}}\leq 0.953$, $\sigma_{4}^{\textup{MC}}\leq 0.916$, $\sigma_{3}^{\textup{IS}}\leq 0.476$, and $\sigma_{4}^{\textup{IS}}\leq 0.457$.
We prove this result in Appendix A.2. For reference, we have collected our results and how they compare to previous results in Table 1. Notably, we obtain upper and lower bounds on $\sigma_{4}^{\textup{MC}}$, $\sigma_{3}^{\textup{IS}}$, and $\sigma_{4}^{\textup{IS}}$ that match up to a small absolute error of $0.005$, so both Theorems 3.1 and 3.2 are nearly optimal for these cases. In particular, this raises the exciting possibility that one might be able to obtain clues toward complete characterizations of $\sigma_{d}^{\textup{MC}}$ and $\sigma_{d}^{\textup{IS}}$ by studying the proofs of Theorems 3.1 and 3.2.
LB ([26]) LB (Theorem 3.1) UB (Theorem 3.2) UB ([23, 18])
$\textup{MAX-CUT},d=3$ 0.944 0.944 0.953 0.971
$\textup{MAX-CUT},d=4$ 0.875 0.911 0.916 0.933
$\textup{MAX-IND-SET},d=3$ 0.458 0.472 0.476 0.485
$\textup{MAX-IND-SET},d=4$ 0.428 0.454 0.457 0.464
Table 1: Comparison of hardness (LB) and algorithms (UB) for certifying upper bounds on MAX-CUT and MAX-IND-SET for $G(n,d)$, in terms of lower bounds and upper bounds on $\sigma_{d}^{\textup{MC}}$ or $\sigma_{d}^{\textup{IS}}$ (see 3.3). The LB results are conditional on 3.1. Our improvements are highlighted.
Figure 3: 44-regular Ramanujan graph found by AlphaEvolve witnessing the lower bound on $\gamma_{4}^{\textup{MC}}$ in Theorem 3.1.
Comments on finding near-optimal lower bounds with AlphaEvolve.
We conclude this section with some remarks on our methodology in Theorem 3.1, and our results. The constructions given in [26] for $d\in\{3,4\}$ are graphs on up to $10$ vertices that were generated by computer-assisted experimentation. We were able to replicate their results by randomly sampling a large number of regular graphs, albeit with the post-hoc knowledge of the number of self-loops in their construction.
Improving their lower bounds necessitates constructions on many more vertices, as the granularity of $\textup{MC}(G)$ or $\textup{IS}(G)$ is limited by the size of $G$. At this scale, this approach of randomly sampling constructions does not work for two reasons: (1) the space of $d$-regular $n$-vertex graphs blows up combinatorially, meaning random sampling will not find interesting “extremal” graphs, and (2) the complexity of computing $\textup{MC}(G)$ or $\textup{IS}(G)$ is exponential in $n$, so it is hard to even compute the bound for a construction.
As stated before, we used AlphaEvolve as a tool to search for finite Ramanujan graphs witnessing the bounds in Theorem 3.1. We circumvent the above limitation by searching over the space $\mathcal{S}=\bigcup_{n=1}^{\infty}\Omega_{n}\times 2^{[n]}$, where $\Omega_{n}$ is the set of $n$-vertex $d$-regular graphs. That is, we search directly for a pair $(G,S)$, where $G$ is a graph and $S$ is a cut (or independent set) in $G$, scoring the pair as $\texttt{score}(G,S)=$-$\infty$ if $G$ is not Ramanujan, and $\texttt{score}(G,S)={|\delta_{G}(S,\bar{S})|}/{|E(G)|}$ otherwise. Here, $\delta_{G}(S,\bar{S})$ is the set of edges crossing the cut $(S,\bar{S})$ in $G$. This gives an efficiently computable score function that lower bounds $\gamma_{d}^{\textup{MC}}$ (or $\gamma_{d}^{\textup{IS}}$). Consequently, AlphaEvolve isn’t limited to the space of small graphs. Indeed, it outputs constructions on as many as $163$ vertices (for $\gamma_{4}^{\textup{IS}}$).
For $d=3$ for MAX-CUT, AlphaEvolve recovered the existing bound $\gamma_{3}^{\textup{MC}}\geq 17/18$ of [26], but wasn’t able to improve on it. We note that for the sake of conciseness we restricted ourselves to the $d\in\{3,4\}$ cases. We expect our techniques to achieve similar improvements over known bounds on $\gamma_{d}^{\textup{MC}}$ and $\gamma_{d}^{\textup{IS}}$ ([3, 23]) for all reasonably small values of $d>4$ except when d1d-1 is an odd perfect square, when [3] already gives the optimal lower bound.
4 Gadget based NP-Hardness for Approximating MAX-3-CUT
The field of approximation algorithms [39] is concerned with computing rigorous worst-case approximations of functions $f:\Omega\to\mathbb{R}$ that are hard to compute exactly. In particular, in hardness of approximation [1] the main goal is to understand when such approximations are computationally hard, i.e., to prove that even approximately computing $f(x)$ is hard. We work within perhaps the simplest and most well-studied framework of Constraint Satisfaction Problems (CSPs).
Definition 4.1 (CSP).
Let $\Sigma$ be a finite alphabet, and let $P_{1},\ldots,P_{r}$ be a collection of predicates over $\Sigma$, that is, $P_{i}:\Sigma^{\ell_{i}}\to\{0,1\}$ for some $\ell_{i}\in\mathbb{N}$. We define $\textup{CSP}(P_{1},\ldots,P_{r})$ to be the collection of instances $I:\Sigma^{n}\to\mathbb{R}_{\geq 0}$ of the form $I(\mathbf{x})=\sum_{j\in[m]}c_{j}(\mathbf{x})$ where for each $j$, $c_{j}$ is a clause, i.e., it takes the form $c_{j}(\mathbf{x})=P_{i}(\mathbf{x}_{a_{1}},\mathbf{x}_{a_{2}},\ldots,\mathbf{x}_{a_{\ell_{i}}})$ for some $i\in[r]$ and distinct $a_{1},\ldots,a_{\ell_{i}}\in[n]$.
In the Max-CSP setting, the input is a description of an instance $I$ of a particular CSP, and the goal is to compute the maximum number of constraints that can be satisfied. In other words, we are trying to approximately compute the function $f(I)=\max_{\mathbf{x}\in\Sigma^{n}}I(\mathbf{x})$. It will be useful to cast the approximation problem as a decision problem, as we do below by introducing the notion of $(c,s)$-approximation.
Definition 4.2 ($(c,s)$-approximation, $\alpha$-approximation).
Given a constraint satisfaction problem $\textup{CSP}(P_{1},\ldots,P_{r})$ and constants $0\leq s\leq c\leq 1$, the $(c,s)$ approximation problem for $\textup{CSP}(P_{1},\ldots,P_{r})$ is the problem of, given an $n$-variable instance $I=\sum_{j\in[m]}c_{j}$, deciding whether (1) $\max_{x}I(x)\geq c\cdot m$ or (2) $\max_{x}I(x)\leq s\cdot m$. Similarly for $0<\alpha\leq 1$, the $\alpha$-approximation problem is the problem of finding some $x^{\prime}$ such that $I(x^{\prime})\geq\alpha\cdot\max_{x}I(x)$.
In particular, a $s/c$-approximation algorithm for the Max-CSP problem gives a $(c,s)$-approximation algorithm. The main driving question can be stated as “given a CSP, what are all pairs $0\leq s\leq c\leq 1$ such that there is an efficient $(c,s)$ approximation for the CSP?”
By now, we have a strong qualitative understanding of this question for all CSPs under the Unique Games Conjecture [33], and a strong quantitative understanding for a large number of CSPs [19, 7]. However, there are some quite simple CSPs for which very little is known quantitatively. For example, consider $\textup{MAX-}{k}\textup{-CUT}$, the problem of partitioning the vertices of a graph into $k$ sets in a way that maximizes the number of edges crossing the partition. This problem can be written as a CSP over alphabet $\Sigma=\mathbb{Z}_{k}$ as $\textup{MAX-}{k}\textup{-CUT}:=\textup{CSP}(P^{\neq})$ where $P^{\neq}$ is the “inequality” predicate $P^{\neq}(x,y)=\mathbf{1}\{x\neq y\}$. For small values of $k$, there is a large gap between the best known algorithmic results (concretely, $0.836$ and $0.857$-approximations for $k=3$ and $k=4$ respectively [16, 9]) and the best known NP-hardness ($16/17+\varepsilon\approx 0.941$ and $85/86+\varepsilon\approx 0.9883$ for $k=3$ and $k=4$ respectively [2]) [^7] Even under the Unique Games Conjecture, the optimal approximability curve is not known for $k=4$, and for $k=3$ is only known in the low completeness regime [20] (which implies a weak inapproximability of $0.989$)..
Gadget reductions provide a methodical framework to find reductions between problems. Specifically in the context of inapproximability, let $\textup{CSP}_{\text{source}}$ be a problem with known inapproximability, and suppose we would like to prove inapproximability for another problem $\textup{CSP}_{\text{target}}$. The basic idea is to systematize the set of reductions from $\textup{CSP}_{\text{source}}$ to $\textup{CSP}_{\text{target}}$ by considering only “local” reductions that replace each clause of an instance of $\textup{CSP}_{\text{source}}$ with a collection of new variables, along with an instance of $\textup{CSP}_{\text{target}}$, known as a gadget on the variables involved. Crucially, the gadget for a particular predicate of $\textup{CSP}_{\text{source}}$ is a finite instance of $\textup{CSP}_{\text{target}}$ of a fixed constant size.
We use AlphaEvolve to find such gadget reductions from the well-understood CSP $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ for $k\in\{3,4\}$. Here $3\textup{LIN}({k})$ is defined as $\textup{CSP}(P_{0}^{=},P_{1}^{=},\ldots,P_{k-1}^{=})$ for the $3$-ary “linear equation” predicates $P_{i}^{=}(x,y,z)=\mathbf{1}\{x+y+z=i\ (\text{mod }k)\}$.
Theorem 4.1.
For any $\varepsilon>0$, there is a gadget reduction from $(1-O(\varepsilon),1/3+O(\varepsilon))$-approximating $3\textup{LIN}({3})$ to $(57/62-\varepsilon,55/62+\varepsilon)$-approximating $\textup{MAX-}{3}\textup{-CUT}$. As a consequence, it is NP-hard to $55/57+\varepsilon$-approximate $\textup{MAX-}{3}\textup{-CUT}$.
Theorem 4.2.
For any $\varepsilon>0$, there is a gadget reduction from $(1-O(\varepsilon),1/4+O(\varepsilon))$-approximating $3\textup{LIN}({4})$ to $(0.9349-\varepsilon,0.9238+\varepsilon)$-approximating $\textup{MAX-}{4}\textup{-CUT}$. As a consequence, it is NP-hard to $0.987+\varepsilon$-approximate $\textup{MAX-}{4}\textup{-CUT}$.
In particular, we used AlphaEvolve to search for the gadgets used in these results. Since $\textup{MAX-}{k}\textup{-CUT}$ is a symmetric $2$-ary predicate, we can view a gadget, which is an instance of $\textup{MAX-}{k}\textup{-CUT}$, as an undirected graph (possibly containing parallel edges). In Figures 4 and 5, we provide a visual representation of these instances. We note that our gadgets look extremely different in the $k=3$ and $k=4$ cases; in particular, in the $k=3$ case our gadgets contain at most $3$ parallel copies of each edge, while in the $k=4$ case the gadgets contain as many as $1429$ parallel copies of some edges. As a result, Theorem 4.2 does not have an inapproximability factor that can be written as a rational number with a small denominator.
(a) Gadget reducing $P_{0}^{=}$ to $P^{\neq}$. The nodes $\{1,2,3\}$ are primary variables, and the rest are auxiliary variables. All edges represent single copies of $P^{\neq}$ clauses. The global variables are not pictured as they are unused.
(b) Gadget reducing $P_{1}^{=}$ to $P^{\neq}$. The nodes $\{1,2,3\}$ are primary variables, $\{3,4,5\}$ are global variables, and the rest are auxiliary variables. The edges represent one, two, and three copies of the corresponding $P^{\neq}$ clause, depending on the thickness.
Figure 4: Gadgets found by AlphaEvolve for reducing $3\textup{LIN}({3})$ to $\textup{MAX-}{3}\textup{-CUT}$ (see Appendix B for a more explicit description via edge lists). Note that an edge $(i,j)$ in the graph corresponds to the predicate $P^{\neq}$ applied to variables $i$ and $j$, with thickness proportional to its number of copies in the instance.
Figure 5: Gadget found by AlphaEvolve for reducing $3\textup{LIN}({4})$ to $\textup{MAX-}{4}\textup{-CUT}$. The thickness of an edge $(i,j)$ is proportional to the number of copies of the predicate $P^{\neq}$ applied to variables $i$ and $j$ in the gadget, which is between $1$ and $1429$ (see Appendix B for a more explicit description via edge lists).
Comparison to previous inapproximability results.
To the best of our knowledge, the current best hardness factor for $\textup{MAX-}{4}\textup{-CUT}$ is $85/86+\varepsilon$ [2, Theorem 1.2], which we improve with Theorem 4.2 to $0.987$. The hardness in [2] was obtained by a reduction from $\textup{MAX-}{3}\textup{-CUT}$ to $\textup{MAX-}{k}\textup{-CUT}$, for all $k>3$.
As for $\textup{MAX-}{3}\textup{-CUT}$, we compare our result in Theorem 4.1 to three works [24, 14, 2], who obtained successively stronger NP-hardness results for $\textup{MAX-}{3}\textup{-CUT}$ (albeit in the perfect completeness regime). Our inapproximability ratio of $55/57+\varepsilon$ in Theorem 4.1 beats the $67/68+\varepsilon$ of Kann, Khanna, Lagergren, and Panconesi [^8] This result can be deduced by plugging in the “state of the art” NP-hardness of approximation for MAX-2-CUT [36] into their argument. and $32/33+\varepsilon$ of Guruswami and Sinop [14].
Our result is not strong enough to beat the state of the art of $16/17+\varepsilon$ by Austrin, O’Donnell, Tan, and Wright [2], who use a custom reduction from Label Cover. In contrast, our result does not require any new PCP machinery, utilizing only Håstad’s classic PCP [19]. We note that there is no fundamental barrier we are aware of which limits the gadget based approach from beating the inapproximability ratio of Austrin et al, and it is conceivable that a gadget based reduction from a different source problem to $3\textup{LIN}({3})$ can achieve this.
4.1 Developing a Search Framework for Soundness and Completeness Guarantees
In order to apply AlphaEvolve to this problem, we developed a template for a gadget-based reduction argument and isolated the properties required from the gadget in this argument (see B.1 and Theorem B.2 for details). We scored a candidate gadget by its final performance, that is, the inapproximability ratio that is implied by applying Theorem B.2 to the candidate gadget.
This template requires $k$ separate gadgets $\{I_{i}:0\leq i<k\}$, that correspond to the predicates $\{P_{i}^{=}:0\leq i<k\}$ of $3\textup{LIN}({k})$. An important feature of $\textup{MAX-}{k}\textup{-CUT}$ is its $S_{k}$-symmetry – it is invariant under permuting the alphabet $\Sigma=\mathbb{Z}_{k}$. Unfortunately this feature is not shared by $3\textup{LIN}({k})$, and as a result, our reductions require a way to break this symmetry of $\textup{MAX-}{k}\textup{-CUT}$. We achieve this by adding $k$ new global variables to the standard systematization of gadgets [36].
Finding gadgets for $\textup{MAX-}{3}\textup{-CUT}$.
AlphaEvolve found a good $I_{0}$ gadget on $9$ variables (Figure 3) quite quickly, mainly because the symmetries of $P_{0}^{=}$ work well with $\textup{MAX-}{k}\textup{-CUT}$ when $k=3$. In particular, this gadget did not require the aforementioned global variables. Building on ideas from the TSSW framework [36] we can show that this is the optimal $I_{0}$ gadget in terms of the final inapproximability ratio.
In contrast, finding good $I_{1}$ and $I_{2}$ gadgets required significantly more effort and some new ideas. Unlike $I_{0}$, we found that increasing the number of auxiliary variables consistently led to better performing gadgets, with the only bottleneck being the runtime of computing the performance of a candidate gadget (we comment more on this point in Section 4.2).
The $I_{1}$ gadget we use in the proof of Theorem 4.1 (Figure 4(b)) requires $14$ total variables. We have some limited evidence that our $I_{1}$ gadget is “locally” optimal [^9] This experiment was done with Krishnamurthy (Dj) Dvijotham, Google DeepMind. in the sense that it gives the optimal inapproximability result among all gadgets on $14$ variables with the same witness as $I_{1}$ for all but one satisfying assignment $x$ of $P_{1}^{\neq}$ (see the completeness case of B.1 for what is meant by a “witness”).
All gadgets found for $\textup{MAX-}{3}\textup{-CUT}$ had between 0 and 3 copies of each edge. This is expected, as optimal gadgets found in many previous results (e.g., [36, 21]) contain small integer weights. As a result, we get a clean inapproximability ratio of $55/57$.
Finding gadgets for $\textup{MAX-}{4}\textup{-CUT}$.
The main difference in our process of finding gadgets for $\textup{MAX-}{4}\textup{-CUT}$ was that we had to search over gadgets with many more variables (as many as $19$), so we required an extremely well-performing implementation of the verifier (again, we elaborate in Section 4.2). Even with this optimized verifier, evaluations were quite slow, requiring on the order of one second to evaluate a single gadget on $19$ variables. As a result, it took a lot longer for AlphaEvolve to find a search algorithm that, given query access to this optimized verifier, could find any nontrivial gadget. The final search algorithm AlphaEvolve produced had the distinctive property of searching over weighted gadgets with real-valued weights, resulting in gadgets with a wide variety of weights between $1$ and $1429$, after appropriate scaling and rounding.
4.2 Faster Verification via AlphaEvolve to Explore Larger Gadgets
The main challenge with finding large gadgets is that the cost of scoring a gadget scales exponentially in the number of variables; computing the completeness and soundness parameters described in B.1 essentially amount to solving an instance of $\textup{MAX-}{k}\textup{-CUT}$, which requires exponential time. Even for $k=3$, AlphaEvolve slows down significantly when searching for gadgets of size as few as $11$ with a brute force $\textup{MAX-}{3}\textup{-CUT}$ solver.
This problem does not have an off-the-shelf solution in the form of existing fast solvers; it is unlikely that existing solvers [11, 29] can be repurposed to solve $\textup{MAX-}{k}\textup{-CUT}$. To solve this issue, we used AlphaEvolve itself to speed up a naive brute force implementation of $\textup{MAX-}{k}\textup{-CUT}$, scoring a candidate implementation by runtime and correctness. In order to calculate the runtime, we created a synthetic dataset of a wide variety of $\textup{MAX-}{k}\textup{-CUT}$ instances, drawn from $20$ random models with varying amounts of planted structure. We then tasked AlphaEvolve to maximize the number of variables $m$ for which the solver requires at most one second on average to solve instances of size $m$ from our dataset.
The biggest challenge was ensuring that AlphaEvolve does not cheat and find an incorrect solver that is much faster. As mentioned before, we achieved correctness by (1) checking that the solver is correct on our synthetic dataset, and (2) using a separate judge LLM to certify that a candidate solver is correct. Both of these techniques were individually too lenient to avoid incorrect solvers, but we found by human inspection that they were enough in combination.
We note that once $m$ is large enough, it is not possible to label our dataset with the “ground truth” scores using a brute-force implementation (which is guaranteed to be correct). Instead, we inductively rely on the correctness of previous solvers produced by AlphaEvolve (which have already passed the above correctness checks) that are fast enough to provide labels for large $m$.
We remark that we did this separately for the $k=3$ and $k=4$ cases, obtaining solvers that are optimized to each particular problem. We found that systems-level improvements were most important to the $k=3$ case, with the final solver offloading the main $O(3^{m})$ time computation to a highly optimized tensor contraction operation in numpy, and only performing slow python-based computation for $O(3^{m/3})$ time. This solver resulted in a $10,000\times$ speedup for instances of size $m=14$.
As for the $k=4$ case, the final solver used a more sequential branch-and-bound strategy along with some systems-level improvements. At $m=19$, this provided a $10,000\times$ speedup against even a numba-accelerated brute force solver.
4.3 Comparison to other computational techniques
We now survey some other computational techniques to find gadgets and discuss why they appear to be infeasible at the scale of our specific problems, even for the simpler setting of $\textup{MAX-}{3}\textup{-CUT}$.
The most straightforward method is the TSSW framework [36], which casts the task of finding the optimal gadget as a linear program. The main difficulty in doing this is the presence of existential quantifiers in computing the completeness of the gadget. In order to eliminate these quantifiers, [36] canonicalize the auxiliary variables in the gadget. As a result, the size of the LP encoding the optimal gadget is doubly exponential in the number of satisfying assignments of the source predicate; this is $3^{3^{6}}$ for a reduction from $3\textup{LIN}({3})$ to $\textup{MAX-}{3}\textup{-CUT}$, which is computationally infeasible. Sometimes (as is the case for our $I_{0}$ gadget for $\textup{MAX-}{3}\textup{-CUT}$) it is possible to argue that not all auxiliary variables are required, leading to a more tractable LP, but it is not clear that this is possible outside of very special cases.
If one wants to fix the number of variables in the gadget to a particular constant smaller than $3^{3^{6}}$ (like $14$ for our gadgets), it is possible to write a mixed integer program (MIP) instead of an LP by encoding the existential constraints using integer variables. For example, even solving the MIP with all but one existential constraint eliminated took $10$ hours [^10] This experiment was done with Krishnamurthy (Dj) Dvijotham, Google DeepMind.. (We used the SCIP solver, with a direct encoding of the problem.) Consequently the MIP approach also seems infeasible with state-of-the-art solvers.
5 Discussion on AI-assisted Mathematics and Complexity Theory
We begin this section with a discussion of the norms by which to understand how AI can assist research in the mathematical sciences. We then discuss the role of AlphaEvolve in achieving the results in this paper, highlighting how our work differs from prior work in computer-assisted theorem proving.
In studying the role of AI in assisting mathematical discovery, we must consider at least four scenarios:
  1. 1.
    A human invokes a language model to summarize the state of prior art, to chart a research plan towards new theorems, or to directly generate fragments of (or entire) proofs.
  2. 2.
    A human uses AI-derived tools such as AlphaEvolve to generate better proof elements (gadgets, graphs).
  3. 3.
    A human uses custom code independent of AI to discover better proof elements.
  4. 4.
    A human discovers the same or better proof elements by hand.
The idea that LLMs can usefully generate summaries of prior literature in a field has been mooted about for some time [38, 40, 22, 15], and indeed our experience confirmed this on a number of prompts. The results (modulo occasional hallucination) are a good starting point for deeper exploration and understanding. Interestingly, across many examples we could rapidly generate an overview of the art in an unfamiliar field; we believe this capability will increasingly be used by scientists working across fields so that — for instance — an algebraist can come up to speed with Ramsey Theory. We believe that in time, this will lead to more fluent cross-pollination across disciplines. We have not yet been able to prompt an LLM into providing a usable research plan to obtain new results, and this is the domain of deeper efforts in this space e.g., Google’s Co-Scientist [17] program.
As mentioned in Section 1, there have been prior efforts to generate proofs for open mathematical statements via prompting an LLM directly [32, 6, 31, 8]. However, this approach has seen limited success to date. In some cases an LLM was indeed able to generate the complete and correct proof of a previously unproven statement [6] (although it is possible that a persistent human could perhaps have derived this), but in many cases [32, 31, 8] the LLM could only generate a proof sketch, which ultimately had to be filled in by humans. In general, these proofs/sketches currently require a human to verify correctness. Indeed, our natural attempts to prompt a standard LLM into directly generating the kinds of combinatorial structures in this paper met with failure  [^11] For example, GPT5 and Gemini 2.5 Pro could not generate the $16/17$-NP-hardness gadget for MAX-CUT, even though the gadget is explicitly mentioned in [36, Figure 4.1] — the prompt being Can you generate the gadget that gives the construction for $16/17$ approximation of MAX-CUT under NP hardness?. While this may become feasible as LLM reasoning capabilities improve [27], a formal comparison is beyond the scope of this paper. It is also worth re-emphasizing that the proof constructions generated by AI in this paper always come with a certificate of correctness which can be verified computationally. So, as long as the verification code is sound, there is no further need for human scrutiny.
Next we discuss whether a human could have discovered the combinatorial structures in our proofs by hand. It is conceivable that a human expert or a highly customized computational solver could eventually find these solutions. However, their discovery is likely beyond the reach of simple ”pencil-and-paper” methods, and standard SMT/MIP solvers failed to produce them.
In both the problems we study, computer assisted methods [36, 19, 26] have been used previously. Kunisky and Yu [26] used computational methods to generate $d$-regular Ramanujan graphs (for $d\in\{3,4\}$) with large cut values (or independent sets). While they do not specify their computational approach, we could replicate their results by random sampling of $d$-regular graphs, and testing for the properties by brute force. For reasons mentioned in Section 3[26] could demonstrate their lower bounds for $n\leq 8$, whereas the graphs we find go up to $n=163$.
For the NP-hardness gadget reduction in Section 4, both the soundness and the completeness constraints can be translated into a linear program via skolemization [36]. However, the size of such a linear program is doubly exponential in the number of nodes in the constraint graph. As a result, even for $\textup{MAX-}{3}\textup{-CUT}$, running a linear program (LP) with the canonical number of variables [36] $\approx 3^{3^{6}}$ becomes infeasible with the standard LP solvers we know of. One can directly attack the non-convex program using MIP/SMT solvers, but with the number of constraints being exponential in the number of variables $n\geq 14$, they too did not seem to scale. (See Section 4.3 for a detailed discussion.)
The best test of robustness for AI-assisted mathematical discovery is durability: whether or not over time, new results such as ours are superseded by humans (possibly with computer assistance). We note that some results in AI-assisted mathematics have in fact been matched or improved quickly without the use of AI [13, 5].
Concluding remarks: While our experience here is limited and far from definitive, we believe some early themes are emerging.
First, AI models with agentic feedback can generate intricate mathematical constructions, sometimes exhibiting signs of reasoning. We observed this when our model produced reasoning logic in $\mathbb{Z}_{3}$ to find the gadget for reducing $x+y+z=0\mod 3$ to $x-y\neq 0\mod 3$. (See Figure 6.)
Second, language models can generate research plans and summarize the state of the art [17]. While we have not succeeded in deriving novel results from this, the capability allows non-specialists to quickly learn new domains, which we anticipate will foster greater scientific cross-pollination.
Third, we expect a growing number of proofs of the form “AI got there first” without clear evidence of the form “this couldn’t have been done without AI”. In all of these cases (and arguably, across applications of AI to science), we expect verification to be an ongoing bottleneck. We leave it as an open question whether directly prompting an LLM can replicate our findings.
Looking ahead, it is conceivable that advances in LLM reasoning [27, 17, 37] could be coupled with AlphaEvolve, especially in generating the initial code-snippet and more effective problem specific prompting of the LLMs used by AlphaEvolve. It is possible that our methods can be extended to other problems in extremal combinatorics. We leave exploration of those problems as future directions.
def build_triad_gadget(num_primary_vars, weight_params):
"""Constructs a symmetric triad-based gadget.
This construction is specific to 3-LIN-3. It creates a highly structured
graph based on the idea of a "triad" of nodes for each primary variable.
assert num_primary_vars == 3
# 3 primary vars, each in a triad of 3 nodes = 9 nodes
# _DOMAIN - 1 fixed-color nodes
num_triad_nodes = num_primary_vars * 3
num_fixed_nodes = _DOMAIN - 1
num_nodes = num_triad_nodes + num_fixed_nodes
gadget = nx.Graph()
gadget.add_nodes_from(range(num_nodes))
# Node mapping:
# p_i -> i (primary node for variable i)
# a_i -> i + num_primary_vars
# b_i -> i + 2 * num_primary_vars
p = list(range(num_primary_vars))
a = [i + num_primary_vars for i in p]
b = [i + 2 * num_primary_vars for i in p]
primary_vars = p
fixed_nodes = list(range(num_triad_nodes, num_nodes))
# 1. Intra-triad edges (to force 3-coloring within a triad)
for i in range(num_primary_vars):
gadget.add_edge(p[i], a[i], weight=weight_params[’W_triad’])
gadget.add_edge(p[i], b[i], weight=weight_params[’W_triad’])
gadget.add_edge(a[i], b[i], weight=weight_params[’W_triad’])
# 2. Inter-triad edges (to enforce the 3-LIN constraint)
for i in range(num_primary_vars):
for j in range(i + 1, num_primary_vars):
gadget.add_edge(p[i], p[j], weight=weight_params[’W_p’])
gadget.add_edge(p[i], a[j], weight=weight_params[’W_pa’])
gadget.add_edge(p[i], b[j], weight=weight_params[’W_pa’])
gadget.add_edge(p[j], a[i], weight=weight_params[’W_pa’])
gadget.add_edge(p[j], b[i], weight=weight_params[’W_pa’])
gadget.add_edge(a[i], a[j], weight=weight_params[’W_aa’])
gadget.add_edge(b[i], b[j], weight=weight_params[’W_aa’])
gadget.add_edge(a[i], b[j], weight=weight_params[’W_ab’])
gadget.add_edge(b[i], a[j], weight=weight_params[’W_ab’])
# 3. Edges to fixed-color nodes (to break symmetries and ground values)
if num_fixed_nodes > 0:
for i in range(num_primary_vars):
for f_node in fixed_nodes:
gadget.add_edge(p[i], f_node, weight=weight_params.get(’W_pf’, 0.0))
gadget.add_edge(a[i], f_node, weight=weight_params.get(’W_af’, 0.0))
gadget.add_edge(b[i], f_node, weight=weight_params.get(’W_af’, 0.0))
return gadget, primary_vars, fixed_nodes
def constraint_graph(num_primary_vars):
Generate the constraint graph using a ’Symmetric Triad Decomposition’ search.
This crazy method abandons the general search over random graphs and instead
proposes a highly symmetric, parameterized structure called a ’triad gadget’.
It then searches for the optimal weight parameters for this structure using
a hill-climbing algorithm with random restarts to escape local optima.
This injects strong structural assumptions into the search, hoping that the
optimal gadget possesses these symmetries.
# Rest of the code follows.
Figure 6: Code generated by AlphaEvolve for reducing $x+y+z=0\mod 3$ to $x-y\neq 0\mod 3$.
6 Acknowledgments
We thank Adam Zsolt Wagner for helping us with AlphaEvolve throughout this project, and for his invaluable advice on our experimental setup — his insights speeded up our experimentation significantly. Swarat Chaudhuri not only helped us with our initial work on MAX-3-CUT, but also generously advised us throughout on various approaches to verification — we are deeply grateful to him for this. Sushant Sachdeva worked closely with us on MAX-3-CUT, and the design of verifiers, and freely shared his valuable intuition on gadget reductions. Pasin Manurangsi replicated some of our early gadgets by hand (especially the optimal gadget that reduces from $P_{0}^{=}$ to $P^{\neq}$), boosting our confidence early on that AlphaEvove was progressing in the right direction. Sidhanth Mohanty made critical contributions in improving the upper bounds on average-case hardness (Theorem 3.2). Krishnamurthy (Dj) Dvijotham helped us with the advanced use of MIP solvers, and ablation studies with some of our verifiers. His ablation studies convinced us that direct usage of MIP solvers would not scale to the gadget sizes we are operating with. Shuang Song helped us with numerous engineering challenges. We thank Raluca Ada Popa, Mary Chesus, Jonathan Katz, Ravi Kumar, James Manyika, Yossi Matias, Jelani Nelson, Rina Panigrahy, Amit Sahai, Thomas Steinke and Jalaj Upadhyay for their valuable feedback on the manuscript. We thank Uri Feige for pointing out an error in a previous version of Appendix A.2. We thank Four Flynn and Pushmeet Kohli for their continued support and feedback through the course of this project.
We especially thank Venkat Guruswami and Madhu Sudan for their ongoing encouragement for this work, and for their incisive perspectives on the current status of various results in inapproximability. In particular, Venkat helped us with nuances of the state of the art for $\textup{MAX-}{k}\textup{-CUT}$, and pointed us to [14, 2].
Appendix A Proofs of Average Case Theorems
In this section, we prove Theorems 3.1 and 3.2.
A.1 Proof of Theorem 3.1
Below we specify three Ramanujan graphs $G_{4}^{\textup{MC}}$, $G_{3}^{\textup{IS}}$, and $G_{4}^{\textup{IS}}$ in the sparse6 format [28]. We also specify a cut $S_{4}^{\textup{MC}}$ (in the form of a subset of vertices), and independent sets $S_{3}^{\textup{IS}}$ and $S_{4}^{\textup{IS}}$ that witness lower bounds on the actual values of $\textup{MC}(G_{4}^{\textup{MC}})$, $\textup{IS}(G_{3}^{\textup{IS}})$, and $\textup{IS}(G_{4}^{\textup{IS}})$. The vertices of all graphs are 0-indexed, as dictated by the sparse6 format. We start by defining $G_{4}^{\textup{MC}}$ and $S_{4}^{\textup{MC}}$:
[…]
0 2 6 7 13 15 16 17 19 20 21 24 25 27 29 31 33 35 36 37 40 42 47 49 50 54 55 56 58 60 61 62 63 65 67 69 72 73 74 77 78 79 83 84 85 86 89 96 99 101 102 103 104 105 110 111 112 113 114 117 120 121 123
$G_{3}^{\textup{IS}}$ and $S_{3}^{\textup{IS}}$ are defined by the following.
[…]
0 15 29 6 4 14 20 28 13 7 12 35 16 30 31 22 32
Finally, $G_{4}^{\textup{IS}}$ and $S_{4}^{\textup{IS}}$ are defined by the following.
[…]
0 1 3 4 7 8 9 10 11 12 13 15 18 24 26 27 28 31 34 40 42 43 44 46 48 53 54 55 56 59 60 61 62 65 75 77 78 82 84 90 91 93 94 98 100 101 103 105 107 109 110 112 114 123 124 125 127 128 129 132 134 135 136 138 141 144 145 146 150 153 154 159 160 162
The following properties can easily be verified and immediately imply Theorem 3.1.
  1. 1.
    $G_{3}^{\textup{IS}}$ is a $3$-regular Ramanujan graph on $36$ vertices. $G_{4}^{\textup{MC}}$ and $G_{4}^{\textup{IS}}$ are $4$-regular Ramanujan graphs on $124$ and $163$ vertices respectively.
  2. 2.
    $S_{3}^{\textup{IS}}$ and $S_{4}^{\textup{IS}}$ are independent sets in $G_{3}^{\textup{IS}}$ and $G_{4}^{\textup{IS}}$ respectively with $|S_{3}^{\textup{IS}}|=17$ and $|S_{4}^{\textup{IS}}|=74$. Consequently, $\textup{IS}(G_{3}^{\textup{IS}})\geq 17/36$, and $\textup{IS}(G_{4}^{\textup{IS}})\geq 74/163$.
  3. 3.
    The number of edges in $G_{4}^{\textup{MC}}$ with exactly one endpoint in $S_{4}^{\textup{MC}}$ is equal to $226$. Consequently, $\textup{MC}(G_{4}^{\textup{MC}})\geq 113/124$.
It can be easily verified that all the graphs are Ramanujan proving the claimed lower bounds on $\gamma_{4}^{\textup{MC}}$, $\gamma_{3}^{\textup{IS}}$, and $\gamma_{4}^{\textup{IS}}$.
A.2 Proof of Theorem 3.2
The goal of this section is to prove improved upper bounds on $\sigma_{d}^{\textup{MC}}$ and $\sigma_{d}^{\textup{IS}}$. The same upper bounds apply to $\gamma_{d}^{\textup{MC}}$ and $\gamma_{d}^{\textup{IS}}$, as will be clear in the proof. We will give a detailed argument for the Max-Cut case, and sketch the (minor) changes required for Max-Independent-Set at the end.
As with previous refutation algorithms [23, 18], our efficient certificate will be a bound on the spectral expansion of a random graph $G\sim G(n,d)$, which is defined as $\lambda^{*}(G)=\max_{i>1}|\lambda_{i}|$ where $\lambda_{1}\geq\lambda_{2}\geq\ldots\geq\lambda_{n}$ are the eigenvalues of the adjacency matrix of $G$. Friedman’s theorem proves that $\lambda^{*}$ concentrates around $2\sqrt{d-1}$.
Theorem A.1 ([12]).
For all $d\in\mathbb{N}$ $\varepsilon>0$, with probability $1-o_{n}(1)$ over $G\sim G(n,d)$, we have $\lambda^{*}(G)\leq 2\sqrt{d-1}+\varepsilon$.
Definition A.1 ($d$-ary tree).
The $d$-ary tree of depth $L$ $T_{d,L}$ is a rooted tree, where the root has $d$ children, and all non-root vertices at distance <L<L from the root have d1d-1 children.
Our improvement on Hoffman’s classical bounds consists of concluding stronger bounds on $\textup{MC}(G)$ and $\textup{IS}(G)$ given that $\lambda^{*}(G)\leq\lambda$ for $\lambda=2\sqrt{d-1}+\varepsilon$. Let us denote by $\Omega_{d,\lambda}$ the set of $d$-regular graphs $G$ such that $\lambda^{*}(G)\leq\lambda$. We will require the notion of a labeled $d$-ary tree, which we define below.
Definition A.2 (Labeled $d$-ary tree).
A $\{\pm 1\}$-labeling of $T_{d,L}$ is a mapping $y$ from the vertices of $T_{d,L}$ to $\{\pm 1\}$.
Let $\Lambda_{d,L}$ be the collection of distributions over $\{\pm 1\}$-labelings of $T_{d,L}$, where the vertices are labeled by elements of $\{\pm 1\}$. For most of the remainder of this section, we will show how we obtain our upper bounds on $\textup{MC}(G)$ given $G\in\Omega_{d,\lambda}$. Our upper bound will be parametrized by a nonnegative integer $L\in\mathbb{N}$. Interestingly, we will exactly recover Hoffman’s bound [23] when $L=1$. The idea is essentially to write a finite LP relaxation of the problem of finding the supremum of the cut fraction ${|\delta_{G}(S,\bar{S})|}/{|E(G)|}$ over all $G\in\Omega_{d,\lambda}$, by projecting all the information about $G$ and $S$ onto “local” neighborhoods at distance $L$.
Let $G\in\Omega_{d,\lambda}$ be an $n$-vertex graph, and let $S\subseteq V(G)$ be a cut in $G$. Let $\alpha\in[0,1/2]$ be a value that approximates $|S|/n$, in the sense that $|\alpha-|S|/n|\leq\delta$. Define the cut vector $x\in\{-1,1\}^{n}$ by $x_{i}=(-1)^{\mathbf{1}\{i\in S\}}$. We will associate the pair $(G,S)$ with a distribution $\mu_{G,S}\in\Lambda_{d,L}$, which can be sampled from as follows:
  • Sample a random ordering of the neighboring edges of each vertex in $G$.
  • Sample a random vertex $i\sim[n]$.
  • We will label $T_{d,L}$ by the values of $x$ at all length $\leq L$ nonbacktracking walks from $i$ (that is, walks that don’t use the same edge twice in a row). Formally, a vertex $v$ of $T_{d,L}$ can be written as a sequence $(q_{1},\ldots,q_{k})$ for $0\leq k\leq L$, $1\leq q_{1}\leq d$, and $1\leq q_{l}\leq d-1$ for $l>1$. We will label $v$ by the $x$ value of the final vertex in the corresponding path starting from $i$ in $G$. In particular, let $i_{0}=i$, and for $1\leq l\leq k$ let $i_{l}$ be the $q_{l}^{th}$ neighbor of $i_{l-1}$ in the sampled ordering, excluding $i_{l-2}$ if $l>1$. We will label $v$ by $y_{v}:=x_{i_{k}}$.
  • Output the resulting labeling $y$ of $T_{d,L}$.
We will deduce some linear constraints on $\mu_{G,S}$, in terms of its pdf.
Lemma A.1 (Local Consistency Constraints).
Let $r_{0}$ be the root of $T_{d,L}$, and let $r_{1}$ be a child of $r_{0}$. For $i\in\{0,1\}$, let $T_{i}$ be the subtree of $T_{d,L}$ rooted at $r_{i}$ containing all vertices at distance at most $L$ from $r_{1-i}$. Then, $T_{0}$ is isomorphic to $T_{1}$, and the marginal distribution of $\mu_{G,S}$ on $T_{0}$ is identical (up to the isomorphism) of that on $T_{1}$.
Proof.
The marginal distribution on $T_{0}$ can be viewed as the following: sample a random vertex $i$ in $G$ and a random neighbor $j\sim i$. Output the labels $x$ at all endpoints of a nonbacktracking walk from $i$ of length at most $L$ if $(i,j)$ is the first edge in the walk, and length at most L1L-1 otherwise.
The marginal distribution on $T_{1}$ is identical, with the roles of $i$ and $j$ switched. Since the distribution of $(i,j)$ is identical to that of $(j,i)$, both marginal distributions are identical. ∎
Lemma A.2 (Average Value of root).
We have $2\alpha-1-2\delta\leq\mathbb{E}_{y\sim\mu_{G,S}}[y_{0}]\leq 2\alpha-1+2\delta$.
Proof.
We can directly compute $\mathbb{E}_{y\sim\mu_{G,S}}[y_{0}]=\mathbb{E}_{i\sim[n]}[x_{i}]=2|S|/n-1$. The sandwiching inequalities follow from the fact that $|\alpha-|S|/n|\leq\delta$. ∎
Lemma A.3 (Spectral Constraints).
The bound $\lambda^{*}(G)\leq\lambda$ implies the bounds
$(2\alpha-1-2\delta)^{2}\leq\mathbb{E}_{y\sim\mu_{G,S},\ell}[y_{0}\cdot y_{\ell}]\leq\left(2\alpha-1+2\delta\right)+(\lambda/d)^{L}$
if $L$ is even, and
$(2\alpha-1-2\delta)^{2}-(\lambda/d)^{L}\leq\mathbb{E}_{y\sim\mu_{G,S},\ell}[y_{0}\cdot y_{\ell}]\leq\left(2\alpha-1+2\delta\right)^{2}+(\lambda/d)^{L}$
if $L$ is odd. Here 0 denotes the root, and $\ell$ denotes the (random) endpoint of an $L$-step random walk in $T_{d,L}$ starting from the root.
Proof.
We begin by computing
$\mathbb{E}_{y\sim\mu_{G,S},\ell}[y_{0}\cdot y_{\ell}]=\frac{x^{T}(A/d)^{L}x}{n}=\frac{x^{T}A^{L}x}{d^{L}\cdot n}.$
Let us write $A=d\frac{\mathbf{1}\mathbf{1}^{T}}{n}+\widetilde{A}$, where $d=\lambda_{1}$ is the trivial eigenvalue of $A$ and $\widetilde{A}$ is the portion of $A$ orthogonal to the trivial eigenvector $\mathbf{1}$. That is, $\widetilde{A}\mathbf{1}=\mathbf{0}$. Note that $\|\widetilde{A}\|_{op}\leq\lambda$. Let us proceed with our computation.
$\mathbb{E}_{y\sim\mu_{G,S},\ell}[y_{0}\cdot y_{\ell}]$ $=\frac{d^{L}\cdot x^{T}\mathbf{1}\mathbf{1}^{T}x}{d^{L}\cdot n^{2}}+\frac{x^{T}\widetilde{A}^{L}x}{d^{L}\cdot n}$
$=\frac{(2|S|-n)^{2}}{n^{2}}+\frac{x^{T}\widetilde{A}^{L}x}{d^{L}\cdot n}.$
The Lemma follows from (1) the fact that $|\alpha-|S|/n|\leq\delta$, and (2) the bounds $|x^{T}\widetilde{A}^{L}x|\leq\|\widetilde{A}\|_{op}\cdot\|x\|_{2}^{2}\leq\lambda n$, along with the bound $x^{T}\widetilde{A}^{L}x\geq 0$ if $L$ is even. ∎
Lemma A.4 (Objective Value).
The cut fraction ${|\delta_{G}(S,\bar{S})|}/{|E(G)|}$ equals $(1-\mathbb{E}_{y\sim\mu_{G,S},i\sim[d]}[y_{0}\cdot y_{i}])/{2}$, where 0 denotes the root, and $i$ denotes the $i^{th}$ neighbor of the root of the $d$-regular tree of depth $L$.
Proof.
This is again a direct computation. We can write
$\mathbb{E}_{y\sim\mu_{G,S},i\sim[d]}[y_{0}\cdot y_{i}]=\frac{x^{T}Ax}{nd}=1-2\frac{|\delta_{G}(S,\bar{S})|}{|E(G)|},$
and rearranging gives the identity. ∎
Lemmas A.1, A.2 and A.3 are all linear constraints in the pdf of the distribution $\mu=\mu_{G,S}$, and Lemma A.4 is a linear objective function. We may solve this LP in the variables $\mu$ for small values of $L$ and $d$ to obtain valid upper bounds on objective function, the cut value $|\delta_{G}(S,\bar{S})|/|E(G)|$. If we solve this LP for a collection of discrete values $\alpha\in\{0,\delta,2\delta,\ldots,1\}$, we obtain an upper bound on the cut value of any cut of a graph $G\in\Omega_{d,\lambda}$.
Computational Efficiency.
We remark that the above LP appears intractable – in particular, it has $2^{46}$ variables even for $d=3$, $L=4$. We apply two optimizations to make it more tractable in practice.
  1. 1.
    Use the same variable for isomorphic labelings of $T_{d,L}$. In particular, if there is a graph isomorphism of $T_{d,L}$ that maps a labeling $y$ to another labeling $y^{\prime}$, then Lemma A.1 implies that they should receive the same probability in $\mu$.
  2. 2.
    Only consider “locally optimal” labelings of $T_{d,L}$. To do this, we apply a preprocessing step, where we ensure that the cut $S$ is locally optimal in the sense that for any vertex $i\in S$, the labeling of $T_{d,L}$ generated by restricting $x\in\{\pm 1\}^{n}$ to the $L$-step neighborhood of $i$ corresponds to a maximum cut of $T_{d,L}$, subject to the “boundary conditions” of labels of the leaves of $T_{d,L}$. Therefore, we can also discard any variables corresponding to such suboptimal labelings of $T_{d,L}$.
Recall that $\lambda=2\sqrt{d-1}+\varepsilon$ for arbitrarily small $\varepsilon$. We will now pick $\varepsilon=10^{-5}$. Using the above optimizations, we are able to solve the $1/\delta$ LPs for $d=3,L=4$ (with $4396$ variables), and $d=4,L=2$ (with $88$ variables) for $\delta=0.0005$, resulting in the universal upper bounds $\textup{MC}(G)\leq 0.953$ for $G\in\Omega_{3,\lambda}$ and $\textup{MC}(G)\leq 0.916$ for $G\in\Omega_{4,\lambda}$.
An interesting point is that the $d=4,L=3$ case has $11880$ variables and is technically within reach of current solvers, but resulted in the same bound on $\sigma^{\textup{MC}}_{4}$ as the $d=4,L=2$ case.
Upper bounds on $\textup{IS}(G)$.
We conclude with some comments on the modifications needed to obtain upper bounds on $\sigma_{d}^{\textup{IS}}$. The main modification is that we need to restrict ourselves to distributions in $\Lambda_{d,L}$ that are supported only on $\{\pm 1\}$-labelings that correspond to actual Independent Sets of $T_{d,L}$. That is, we only consider labelings $y$ such that for any edge $\{i,j\}$ in $T_{d,L}$, $y_{i}$ and $y_{j}$ are not both equal to $1$. The objective function disappears, and we now only need to find the largest possible value $\alpha$ such that the LP is still feasible. The $1/\delta$ LPs are tractable for $d=3,L=4$ (with $1771$ variables) and $d=4,L=3$ (with $8855$ variables), and they result in the bounds $\sigma_{3}^{\textup{IS}}\leq 0.476$ and $\sigma_{4}^{\textup{IS}}\leq 0.457$.
Appendix B Proofs of Theorems 4.1 and 4.2
In this section, we give the details of the NP-hardness of approximation proofs for MAX-k-CUT for $k\in\{3,4\}$. Let us begin by formally defining a gadget, specifically in the context of the reduction from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$. We will need $k$ separate gadgets. In particular, for each defining predicate $P_{i}^{=}$ of $3\textup{LIN}({k})$, we will need an $i$-gadget, which we define below.
Definition B.1 (Gadgets).
Let $k\geq 2$ and define $\Sigma=\mathbb{Z}_{k}$.
  • $i$-gadget: For $i\in\Sigma$, an $i$-gadget from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ is an instance $I_{i}$ of $\textup{MAX-}{k}\textup{-CUT}$ over $3+k+n_{\textup{aux}}$ variables, where $n_{\textup{aux}}\in\mathbb{N}$.
  • Primary, global, and auxiliary variables: We partition the set of variables $3+k+n_{\textup{aux}}]$ into three primary variables $\{1,2,3\}$, $k$ global variables $\{4,5,\ldots,k+3\}$, and $n_{\textup{aux}}$ auxiliary variables $\{k+4,\ldots,3+k+n_{\textup{aux}}\}$. We will always write an assignment to the variables as a tuple $(x,y,z)\in\Sigma^{3+k+n_{\textup{aux}}}$ where $x\in\Sigma^{3}$, $y\in\Sigma^{k}$, and $z\in\Sigma^{n_{\textup{aux}}}$.
The variables $y$ will be restricted in a way to break the symmetries of $\textup{MAX-}{k}\textup{-CUT}$. Concretely, let $Y$ be the set of strings $y$ in $\Sigma^{k}$ that are lexicographically not larger than any string $y^{\sigma}=(\sigma(y_{1}),\sigma(y_{2}),\ldots,\sigma(y_{k}))$ that is obtained by permuting the alphabet $\Sigma$ by some permutation $\sigma:\Sigma\to\Sigma$. In particular, note that the vector $(0,1,2,\ldots,k-1)$ is in $Y$. We will only ever assign $y$ to values in $Y$.
Definition B.2 (Soundness and completeness).
We will set $C_{i}=(P_{i}^{=})^{-1}(1)$ and $S_{i}=(P_{i}^{=})^{-1}(1)$ to be the set of satisfying and unsatisfying assignments for the source predicate $P_{i}^{=}$ respectively.
We will associate four real-valued parameters to the gadget $I_{i}=\sum_{j\in[m]}c_{j}$ as follows.
  • (Completeness analysis) $c(I_{i})$ is defined as $\min_{x\in C_{i}}\max_{z\in\Sigma^{n_{\textup{aux}}}}I_{i}(x,y,z)$ for $y=(0,1,\ldots,k-1)$. For $x\in C_{i}$, we say that $\arg\max_{z\in\Sigma^{n_{\textup{aux}}}}I_{i}(x,y,z)$ is the witness for $x$.
  • (Soundness analysis) $c^{\prime}(I_{i})$ is defined as
    $c^{\prime}(I_{i})=\max_{x\in C_{i},z\in\Sigma^{n_{\textup{aux}}},y\in Y}I_{i}(x,y,z).$
    Also, $s(I_{i})$ is defined as
    $s(I_{i})=\max_{x\in S_{i},z\in\Sigma^{n_{\textup{aux}}},y\in Y}I_{i}(x,y,z).$
  • (Number of clauses) $t(I_{i})$ is defined to be $m$, that is, the number of clauses present in $I_{i}$.
We will show that any collection of $i$-gadgets for all $i\in\mathbb{Z}_{k}$ provides a reduction whose performance depends on the various parameters defined above. After this, in Section B.1, we give an explicit description of our gadgets and state the parameters achieved by them.
The starting point of our reduction will be the following Theorem due to Håstad.
Theorem B.1 ([19]).
Let $k\geq 2$. Given an instance $I=\sum_{j\in[m]}c_{j}$ of $3\textup{LIN}({k})$ with $m/k$ clauses from each of the predicates $P_{0}^{=},P_{1}^{=},\ldots,P_{k-1}^{=}$, it is NP-hard to distinguish between the following cases for any $\varepsilon>0$:
  1. 1.
    (Completeness case) $\max_{x}I(x)\geq(1-\varepsilon)\cdot m$, or
  2. 2.
    (Soundness case) $\max_{x}I(x)\leq(1/k+\varepsilon)\cdot m$.
Theorem B.2 details how our gadgets imply NP-hardness of approximation for $\textup{MAX-}{k}\textup{-CUT}$, with $k\in\{3,4\}$.
Theorem B.2.
Let $k\geq 2$, and set $\Sigma=\mathbb{Z}_{k}$. Let $I_{i}$ be an $i$-gadget from $3\textup{LIN}({k})$ to $\textup{MAX-}{k}\textup{-CUT}$ for each $i\in\Sigma$. For any $\varepsilon>0$, given an instance $J$ of $\textup{MAX-}{k}\textup{-CUT}$ with $M$ constraints, it is NP-hard to distinguish between the following two cases for any $\varepsilon>0$:
  1. 1.
    (Completeness case) $\max_{x}J(x)\geq(a-\varepsilon)\cdot M$, or
  2. 2.
    (Soundness case) $\max_{x}J(x)\leq(b+\varepsilon)\cdot M$,
where
$a=\frac{\sum_{i\in\Sigma}c(I_{i})}{\sum_{i\in\Sigma}t(I_{i})},\quad b=\frac{c^{\prime}(I_{r_{0}})+\sum_{i\in\Sigma\setminus\{0\}}s(I_{r_{i}})}{\sum_{i\in\Sigma}t(I_{i})}.$
Here $r_{0},r_{1},\ldots r_{k-1}$ is a permutation of $\{0,1,\ldots,k-1\}$ such that $c^{\prime}(I_{r_{i}})-s(I_{r_{i}})$ is maximized by $r_{0}$.
Proof of Theorem B.2.
We will perform a reduction starting from the NP-hard problem in Theorem B.1 to $\textup{MAX-}{k}\textup{-CUT}$. Let $I=\sum_{j\in[m]}c_{j}$ be an $n$-variable instance of $3\textup{LIN}({k})$, and suppose that it has $m/k$ clauses using $P_{i}^{=}$ for all $i\in\Sigma$. We will show how to reduce $I$ to an instance $J$ of $\textup{MAX-}{k}\textup{-CUT}$.
Let $n_{\textup{aux}}$ be the number of auxiliary variables in all gadgets $I_{i}$; we can assume they all have the same number of auxiliary variables by introducing new dummy variables to some of them. $J$ will be defined as an instance on $n+k+n_{\textup{aux}}\cdot m$ variables, which we will denote as $x\in\Sigma^{n}$, $y\in\Sigma^{k}$, and $z\in\Sigma^{n_{\textup{aux}}\cdot m}$. For $j\in[m]$ we will denote the $j^{th}$ block of size $n_{\textup{aux}}$ in $z$ as $z^{j}$. We will refer to $x,y,z$ as the primary, global, and auxiliary variables respectively.
Let $c_{j}=P^{=}_{i}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$ be a clause corresponding to the predicate $P^{=}_{i}$ for some $i$. We will add a copy of $I_{i}$ on the variables $(x^{\prime},y,z^{j})$, where $x^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$. Note that the number of clauses in $J$ is $M=(m/k)\cdot\sum_{i\in\Sigma}t(I_{i})$.
Now we will invoke Theorem B.1, which says that it is NP-hard to distinguish between the cases that $I$ is almost completely satisfiable, and $I$ is essentially $1/k$-satisfiable.
  1. 1.
    (Completeness case). Assuming there is an $x\in\Sigma^{n}$ such that $I(x)\geq(1-\varepsilon)\cdot m$, we show how to construct an assignment $(x,y,z)\in\Sigma^{n+k+n_{\textup{aux}}\cdot m}$ such that $J(x,y,z)\geq a\cdot M$. We will set $y=(0,1,2,\ldots,k-1)$, and for each clause $c_{j}=P_{i}^{=}(x_{j_{1}},x_{j_{2}},x_{j_{3}})$, $z^{j}$ will be set as the witness $z_{j}=\arg\max_{z^{j}\in\Sigma^{n_{\textup{aux}}}}I_{i}(x^{\prime},y,z^{j})$, where $x^{\prime}=(x_{j_{1}},x_{j_{2}},x_{j_{3}})$.
    It is immediate that $(x,y,z)$ satisfies at least
    $\sum_{i\in\Sigma}c(I_{i})\cdot m/k-O(\varepsilon\cdot m)=a\cdot M-O(\varepsilon)\cdot m\geq(a-\varepsilon^{\prime})\cdot M$
    clauses of $J$, where $\varepsilon^{\prime}=O(\varepsilon)$ can be chosen to be arbitrarily close to 0.
  2. 2.
    (Soundness case). We must show that if $\max_{x}I(x)\leq(1/3+\varepsilon)\cdot m$, then $\max_{x,y,z}J(x,y,z)\leq(b+\varepsilon^{\prime})\cdot M$. Let $(x,y,z)$ be an assignment to the variables of $J$.
    The first step we perform is to assume without loss of generality that $y\in Y$. To do this, note that by definition of $Y$, there is some permutation $\sigma:\Sigma\to\Sigma$ such that $y^{\sigma}=(\sigma(y_{1}),\ldots,\sigma(y_{k}))\in Y$. We will apply $\sigma$ to all variables in the full assignment $(x,y,z)$, resulting in no change in the value of $J(x,y,z)$ while guaranteeing that $y\in Y$.
    For $i\in\Sigma$, let $\alpha_{i}$ denote the fraction of $P^{=}_{i}$ constraints of $I$ that are satisfied by $x$. By assumption, we have $\sum_{i\in\Sigma}\alpha_{i}\leq 1+3\varepsilon$. Let us define $\hat{\alpha_{i}}=\alpha_{i}/(\sum_{i^{\prime}\in\Sigma}\alpha_{i^{\prime}})$, so $\sum_{i\in\Sigma}\hat{\alpha_{i}}=1$. In other words, $\{\hat{\alpha_{i}}:i\in\Sigma\}$ defines a probability distribution.
    On the other hand, since $y\in Y$ we can now bound $J(x,y,z)$ as follows.
    $J(x,y,z)$ $\leq m\cdot\left(\sum_{i\in\Sigma}\alpha_{i}\cdot c^{\prime}(I_{i})+(1-\alpha_{i})\cdot s(I_{i})\right)$
    $\leq m\cdot\left(\sum_{i\in\Sigma}\hat{\alpha_{i}}\cdot c^{\prime}(I_{i})+(1-\hat{\alpha_{i}})\cdot s(I_{i})\right)+O(\varepsilon)\cdot m$ ($\sum_{i\in\Sigma}\alpha_{i}\leq 1+3\varepsilon$)
    $=m\cdot\left(\sum_{i\in\Sigma}\hat{\alpha_{i}}\cdot(c^{\prime}(I_{i})-s(I_{i}))+s(I_{i})\right)+O(\varepsilon)\cdot m.$
    Note that subject to $\hat{\alpha_{i}}$ being a probability distribution, this is maximized when $\hat{\alpha_{i}}$ is a point mass on a maximizer of $c^{\prime}(I_{i})-s(I_{i})$. Therefore, this is bounded by
    $m\cdot\left(c^{\prime}(I_{r_{0}})+\sum_{i\in\Sigma\setminus\{0\}}s(I_{r_{i}})\right)+O(\varepsilon)\cdot m=(b+O(\varepsilon))\cdot M=(b+\varepsilon^{\prime})\cdot M,$
    where $\varepsilon^{\prime}$ can be chosen to be arbitrarily close to 0. This completes the proof of the soundness case.
B.1 Description of our gadgets
For $k\in\{3,4\}$, we used AlphaEvolve to find gadgets $\{I_{i}:i\in\mathbb{Z}_{k}\}$, scoring a particular gadget by the inverse inapproximability ratio $a/b$, where $a$ and $b$ are defined in Theorem B.2. In this part, we concretely define the specific gadgets found by AlphaEvolve and list the parameters they achieve (B.2). We will conclude the final inapproximability results (Theorems 4.1 and 4.2) by applying Theorem B.2 on these gadgets.
We have formatted a gadget as a weighted “edge list” which corresponds directly to a collection of $P^{\neq}$ clauses. The format is meant to be understood as follows: an tuple $(a,b,w)$ corresponds to $w$ parallel edges from variable $a$ to variable $b$, that is, the gadget contains $w$ copies of the term $P^{\neq}$ applied to variables $a$ and $b$.
Gadgets for MAX-3-CUT.
We write the list of edges for $I_{0}$ first, which has $12$ total variables, and hence $6$ auxiliary variables. Note that the global variables $\{4,5,6\}$ are not used in this gadget.
[(1, 7, 1), (1, 8, 1), (1, 9, 1), (1, 10, 1), (2, 9, 1), (2, 10, 1), (2, 11, 1), (2, 12, 1), (3, 7, 1), (3, 8, 1), (3, 11, 1), (3, 12, 1), (7, 12, 1), (7, 10, 1), (8, 9, 1), (8, 11, 1), (9, 12, 1), (10, 11, 1)]
Below, we write the list of edges for our $I_{1}$ gadget, which has $6+n_{\textup{aux}}=14$ total variables, and hence $8$ auxiliary variables. Note that repeated edges correspond to multiple $P^{\neq}$ clauses on the same pair of variables.
[(3, 7, 1), (4, 6, 3), (4, 12, 1), (3, 13, 2), (5, 10, 1), (9, 11, 1), (11, 14, 1), (1, 3, 2), (1, 9, 1), (2, 8, 1), (2, 14, 1), (13, 14, 2), (6, 11, 1), (4, 5, 3), (3, 9, 1), (5, 6, 3), (4, 8, 2), (5, 9, 2), (1, 2, 2), (2, 7, 1), (10, 14, 1), (1, 8, 1), (1, 14, 1), (2, 13, 2), (6, 7, 2), (7, 12, 1), (4, 7, 1), (12, 14, 1), (3, 8, 1), (3, 14, 1), (5, 8, 1), (8, 10, 1), (2, 3, 2), (2, 9, 1), (1, 7, 1), (1, 13, 2), (6, 9, 1)]
$I_{2}$ is defined identically to $I_{1}$, except with the global variables reordered from $(4,5,6)$ to $(4,6,5)$. We can calculate the relevant parameters of these gadgets as below.
Lemma B.1.
Let $I_{0},I_{1}$, and $I_{2}$ be as above. $I_{0}$ is a 0-gadget from $3\textup{LIN}({3})$ to $\textup{MAX-}{3}\textup{-CUT}$ with parameters $c(I_{0})=c^{\prime}(I_{0})=t(I_{0})=18$, and $s(I_{0})=16$. For $i\in\{1,2\}$, $I_{i}$ is an $i$-gadget with parameters $c(I_{i})=c^{\prime}(I_{i})=48$, $s(I_{i})=46$, and $t(I_{i})=53$.
Combined with Theorem B.2, we immediately get Theorem 4.1.
Gadgets for MAX-4-CUT.
Below we write the weighted edge list for our $I_{0}$ gadget for $\textup{MAX-}{4}\textup{-CUT}$.
[(1, 2, 866), (1, 3, 865), (1, 5, 324), (1, 6, 168), (1, 7, 324), (1, 8, 178), (1, 9, 1361), (1, 10, 648), (1, 11, 628), (1, 12, 731), (1, 13, 36), (1, 14, 331), (1, 15, 1038), (1, 16, 473), (1, 17, 73), (1, 18, 442), (1, 19, 1013), (2, 3, 866), (2, 5, 323), (2, 6, 168), (2, 7, 323), (2, 8, 1243), (2, 9, 1361), (2, 10, 724), (2, 11, 261), (2, 12, 731), (2, 13, 36), (2, 14, 331), (2, 15, 65), (2, 16, 218), (2, 17, 601), (2, 18, 463), (2, 19, 995), (3, 5, 324), (3, 6, 168), (3, 7, 322), (3, 8, 1125), (3, 9, 1360), (3, 10, 288), (3, 11, 719), (3, 12, 731), (3, 13, 37), (3, 14, 331), (3, 15, 1037), (3, 16, 509), (3, 17, 603), (3, 18, 149), (3, 19, 66), (4, 5, 1261), (4, 6, 1089), (4, 7, 1259), (4, 9, 173), (4, 10, 19), (4, 11, 11), (4, 12, 167), (4, 13, 1255), (4, 14, 12), (4, 16, 11), (4, 17, 9), (4, 18, 29), (5, 6, 660), (5, 7, 1429), (5, 8, 87), (5, 9, 314), (5, 10, 34), (5, 11, 10), (5, 12, 581), (5, 13, 2), (5, 14, 389), (5, 15, 8), (5, 16, 15), (5, 17, 24), (5, 18, 7), (5, 19, 16), (6, 7, 656), (6, 8, 380), (6, 10, 181), (6, 11, 159), (6, 13, 49), (6, 15, 350), (6, 16, 78), (6, 17, 196), (6, 18, 79), (6, 19, 322), (7, 8, 94), (7, 9, 319), (7, 10, 35), (7, 11, 20), (7, 12, 655), (7, 13, 1), (7, 14, 316), (7, 15, 16), (7, 16, 2), (7, 17, 23), (7, 18, 7), (7, 19, 22), (8, 9, 52), (8, 10, 338), (8, 11, 460), (8, 12, 508), (8, 13, 20), (8, 14, 338), (8, 16, 326), (8, 17, 59), (8, 18, 388), (9, 10, 361), (9, 11, 343), (9, 12, 263), (9, 13, 1265), (9, 14, 60), (9, 15, 25), (9, 16, 288), (9, 17, 199), (9, 18, 207), (9, 19, 25), (10, 11, 3), (10, 12, 344), (10, 13, 226), (10, 14, 97), (10, 15, 404), (10, 16, 3), (10, 17, 3), (10, 18, 103), (10, 19, 82), (11, 12, 296), (11, 13, 360), (11, 14, 160), (11, 15, 60), (11, 16, 109), (11, 17, 2), (11, 19, 436), (12, 13, 898), (12, 15, 517), (12, 16, 120), (12, 17, 264), (12, 18, 238), (12, 19, 448), (13, 14, 344), (13, 15, 34), (13, 16, 262), (13, 17, 300), (13, 18, 178), (13, 19, 6), (14, 15, 190), (14, 16, 218), (14, 17, 134), (14, 18, 66), (14, 19, 242), (15, 16, 97), (15, 17, 328), (15, 18, 352), (16, 17, 2), (16, 19, 252), (17, 19, 357), (18, 19, 73)]
For $i\geq 1$, our $I_{i}$ gadget is obtained from our $I_{i-1}$ gadget by reordering the global variables from $(4,5,6,7)$ to $(7,4,5,6)$. The below Lemma can be verified computationally using Theorem B.1.
Lemma B.2.
Let $I_{0},I_{1}$, $I_{2}$, and $I_{3}$ be as above. For each $i$, $I_{i}$ is an $i$-gadget from $3\textup{LIN}({4})$ to $\textup{MAX-}{4}\textup{-CUT}$ with parameters $c(I_{i})=49535$, $c^{\prime}(I_{i})=49538$, $s(I_{i})=48681$, and $t(I_{i})=52941$.
Together with Theorem B.2, we immediately get Theorem 4.2.
Finally, for the convenience of the reader, we provide a code snippet that computes the set $Y$ (from B.1) of allowed assignments to the global variables as a function of $k$ below.
# Outputs list of allowed assignments to global variables for the 3lin(k) to max-k-cut reduction
def allowed_global_assignments(k):
# iterate over all k-tuples of elements of Z_k
ret = []
for y in itertools.product(range(k), repeat=k):
# To determine whether y is lexicographically minimal with respect to permuting each entry by the same permutation of Z_k, we check that ith distinct element of y is equal to i-1.
elements = []
for element in y:
if element not in elements:
elements.append(element)
if elements == list(range(len(elements))):
ret.append(y)
return ret